OpenStudy (emilyjones284):

Please help ...medal and fan (: A set of data that is normally distributed has a mean of 35.6 and standard deviation of 2.5. Which number is between 1 and 2 standard deviations of the mean?

OpenStudy (nurali):

35.6 + sd = 35.6 + 2.5 = 38.1

OpenStudy (nurali):

35.6 + 2sd = 38.1 + 2.5 = 40.6

OpenStudy (emilyjones284):

So what would the answer be?

jimthompson5910 (jim_thompson5910):

are you given choices?

OpenStudy (nurali):

are you given choices?

OpenStudy (emilyjones284):

Yes, 29, 41.2, 34.1 , or 38.3

jimthompson5910 (jim_thompson5910):

look at the work Nurali posted

OpenStudy (nurali):

Seems like D is between the above 2 values.

jimthompson5910 (jim_thompson5910):

basically this means that the number 38.1 is 1 std dev from the mean and 40.6 is 2 std dev away from the mean so whatever is between those 2 values is your answer

OpenStudy (emilyjones284):

Ohhhhhh okay I get it :D thank you (:

OpenStudy (nurali):

Anytime.